A confined gas is at temperature of 27°C when it has a pressure and volume of 1000 Pascals and 20 m3 and the temperature is raised to 30°C, what is the nevw pressure?
Answer #1

P/T is constant because volume of gas remains fixed.

P/303.15=1000/300.15

temperature is converted into kelvin k=c+273.15

P= 1010 Pa